OpenStudy (anonymous):

Find a quadratic model for the set of values: (-2, -20), (0, -4), (4,-20) Show your work

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (perl):

You need a system of equations

OpenStudy (usukidoll):

I don't feel well today. I woke up with a migraine (again)

OpenStudy (anonymous):

Take some advil and drink water

OpenStudy (perl):

$$ \Large{ y = ax^2 + bx + c \\~\\ -20 = a(-2)^2 + b(-2) + c \\-4 = a(0)^2 + b(0) + c \\ -20 = a(4)^2 + b(4) + c }$$

OpenStudy (anonymous):

i honestly have no clue how to do this equation

OpenStudy (perl):

Do you agree with what I have so far. I plugged in the values into a generic quadratic function.

OpenStudy (anonymous):

I think so, i wouldn't know if you were wrong

OpenStudy (perl):

Finding a quadratic model means fitting the points to a quadratic curve \( \large y = ax^2 + bx + c \) . So our job is to find the parameters \( \large a,b,c\)

OpenStudy (anonymous):

@perl how do i find the a,b,c

OpenStudy (perl):

We have these three equations so far. \[ \\~\\\Large { -20 = a(-2)^2 + b(-2) + c \\-4 = a(0)^2 + b(0) + c \\ -20 = a(4)^2 + b(4) + c }\]

OpenStudy (perl):

the second equation gives us -4 = 0 + 0 + c -4 = c

OpenStudy (anonymous):

So c=-4

OpenStudy (perl):

Yes. Now back substitute.

OpenStudy (perl):

We have these three equations so far. \[ \\~\\\Large { -20 = a(-2)^2 + b(-2) + (-4) \\-4 = a(0)^2 + b(0) + (-4) \\ -20 = a(4)^2 + b(4) + (-4) }\]

OpenStudy (perl):

we don't need the second equation any longer

OpenStudy (perl):

\[ \\~\\\Large { -20 = a(-2)^2 + b(-2) + (-4) \\ -20 = a(4)^2 + b(4) + (-4) }\]

OpenStudy (anonymous):

im writing as we go

OpenStudy (perl):

\[ \\~\\\Large { -20 = a(-2)^2 + b(-2) -4 \\ -20 = a(4)^2 + b(4) -4 \\~\\} \rm simplify \\~\\\Large { -20 = 16a -2b -4 \\ -20 = 16a + 4b -4 \\~\\ \rm simplify ~ more\\~\\ -16 = 16a -2b \\ -16 = 16a + 4b }\]

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (perl):

We can solve for `b`

OpenStudy (anonymous):

So we haven't solved A or C yet?

OpenStudy (perl):

we solved for c = -4

OpenStudy (anonymous):

Ok so how do we solve for A and B than? Im sorry i really dont understand

OpenStudy (anonymous):

@perl

OpenStudy (perl):

we can solve that linear equation

OpenStudy (anonymous):

how?

OpenStudy (perl):

\[ \\~\\\Large { -20 = a(-2)^2 + b(-2) -4 \\ -20 = a(4)^2 + b(4) -4 \\~\\} \rm simplify \\~\\\Large { -20 = 16a -2b -4 \\ -20 = 16a + 4b -4 \\~\\ \rm simplify ~ more\\~\\ -16 = 16a -2b \\ -16 = 16a + 4b \\~\\ \rm add ~ the ~ two ~ equations\\~\\ \\-32 = 2b \\~\\ \\b = \frac{-32}{2} }\]

OpenStudy (anonymous):

So b=-16?

OpenStudy (perl):

yes

OpenStudy (anonymous):

So now we need to find A

OpenStudy (perl):

y = -2x^2 + 4x -4 should be the answer. I have an error somewhere above.

OpenStudy (anonymous):

so b should really be b=4?

OpenStudy (anonymous):

A=-2 B=4 C=-4?

OpenStudy (perl):

yes

OpenStudy (perl):

I have to go back and find the error now ;)

OpenStudy (anonymous):

Awesome thank you so much!

OpenStudy (perl):

but this method is sound if you want to try it again

OpenStudy (perl):

if you ever*

OpenStudy (anonymous):

yeah i prob will, i have a whole other algebra class ahead of me

OpenStudy (perl):

\[ \\~\\\Large { -20 = a(-2)^2 + b(-2) -4 \\ -20 = a(4)^2 + b(4) -4 \\~\\} \rm simplify \\~\\\Large { -20 = \color{red}4a -2b -4 \\ -20 = 16a + 4b -4 \\~\\ \rm simplify ~ more\\~\\ -16 = \color{red}4a -2b \\ -16 = 16a + 4b \\~\\ \rm solve ~this~ system~ of~ equations\\~\\ \\ a = -2 ~ , b = 4 \\~\\ \rm backsubstitute \\y = -2x^2 + 4x - 4 }\]

OpenStudy (perl):

Another way to solve this problem is to note that your parabola has a y intercept of -4, because x=0 at the y intercept. That reduces the problem of solving a linear system of 2 equations as we have above.